# Spatial Scale of Agglomeration and Dispersion: Number, Spacing, and the Spatial Extent of Cities

### v8

@@ 2.1  The Beckmann model: A “local” dispersion force
While $\bar{\bm{x}}$ is an equilibrium, with both agglomeration and dispersion forces, $\bar{\bm{x}}$ is not stable if the former dominates the latter. To assess the stability of $\bar{\bm{x}}$, we examine the incentive of movers. Consider a small mass of agents moving from region 2 to 1, so that $x_{1}$ rises and $x_{2}$ falls symmetrically. If $\Delta$ decreases under this perturbation, then $\Delta<0$ after the move: utility in region 2 exceeds that in region 1, and the movers will prefer to return. This induces a restoring force and hence $\bar{\bm{x}}$ is locally stable. If $\Delta$ increases, then $\Delta>0$ after the move: the movers prefer to stay in region 1, and the deviation attracts additional movers now that the utility is higher in region $1$. Hence, $\bar{\bm{x}}$ is unstable.
@@ 2.2  The Braid model: A “global” dispersion force
Here, $y_{i}(\bm{x})$ is the income of an agent in the region $i$. Each region is endowed with one unit of fixed expenditure, which is allocated to all agents according to accessibility. Suppose that the share received by an agent in region $i$ from region $m$ is given by [^7] Note that the total expenditure of the region $m$ is equal to one: $\sum_{i\in\mathcal{I}}x_{i}s_{i\mid m}=1$. Possible microfoundations for this toy model are abstracted away for brevity.
@@
demonstrating that a marginal increase in agents in any region $j\in\{1,2\}$ has negative impacts on $s_{i\mid m}$. This is simply because it increases the denominator of $s_{i\mid m}$. Whether income $y_{i}=\sum_{j}s_{i\mid j}$ as a whole increases or not after a migration shock depends on the relative magnitudes of these impacts. Specifically, analogous to the utility gain $\omega$, we can compute the income gain of a marginal mover from region $2$ to 11: [^8] From Eq. 9, $\frac{\partial{y_{1}}}{\partial{x_{1}}}=-(s_{1\mid 1})^{2}-(s_{1\mid 2})^{2}=-\frac{1+\phi^{2}}{\bar{x}^{2}(1+\phi)^{2}}$ and $\frac{\partial{y_{2}}}{\partial{x_{1}}}=-s_{2\mid 1}\cdot s_{1\mid 1}-s_{2\mid 2}\cdot s_{1\mid 2}=-\frac{2\phi}{\bar{x}^{2}(1+\phi)^{2}}$.
@@ 3  Many regions
This section examines how the proposed taxonomy of spatial models maps to the endogenous spatial patterns and their comparative statics in an $N$-region economy. All variables and functions (e.g., $\bm{x}$, $\bm{v}$, $\phi_{ij}]$, $\mathbf{D}$) are straightforwardly extended. The set of regions is now denoted by $\mathcal{I}\equiv\{1,2,\ldots,N\}$.
@@
The proximity matrix is given by $\phi_{ij}=\phi^{\ell_{ij}}$, where $\phi\in(0,1)$ is the ease of transport between two consecutive regions, and $\ell_{ij}\equiv\min\{|i-j|,N-|i-j|\}$ is the distance between regions $i$ and $j$ over the circumference. All regions are symmetric regarding their local fundamentals (e.g., innate amenity or productivity). In addition, $N$ is a multiple of four. [^14] This restriction on $N$ is only for expositional simplicity. See Remark 3 in Appendix A.
@@ 3.1  The stability of the symmetric equilibrium
Endogenous agglomeration occurs when the symmetric equilibrium $\bar{\bm{x}}$ is unstable, that is, when it does not withstand small migration shocks. In an economy with $N$ regions, such shocks can be represented as
@@
where $\epsilon$ is a sufficiently small scalar. The vector $\bm{z}=(z_{i})_{i\in\mathcal{I}}$ is a deviation pattern: $z_{i}>0$ indicates an inflow into region $i$ and $z_{i}<0$ an outflow. Since the total population is fixed, we require $\sum_{i\in\mathcal{I}}z_{i}=0$. The $N$-region setting admits a substantially richer set of deviation patterns than the two-region case, because $(\bar{x}+\epsilon,\bar{x}-\epsilon)$ is the only possible perturbation in the two-region setting. Below, we normalize $\|\bm{z}\|=1$.
@@
In the Beckmann model, we again have $\mathbf{V}=-\beta\mathbf{I}+\alpha\mathbf{D}$, where $\mathbf{I}$ and $\mathbf{D}$ are replaced by the $N$-region identity and row-normalized proximity matrices. [^17] Observe that Eq. 2 is defined for general $N$. The $k$th eigenvalue of $\mathbf{V}$ is then $\omega_{k}=-\beta+\alpha\Theta_{k}$ because $\mathbf{D}\bm{z}_{k}=\Theta_{k}\bm{z}_{k}$ implies $\mathbf{V}\bm{z}_{k}=(-\beta+\alpha\Theta_{k})\bm{z}_{k}$. The stability of $\bar{\bm{x}}$ is determined by the sign of $\omega^{*}=\max_{k}\{\omega_{k}\}$. Consider a value of $\phi$ such that $\bar{\bm{x}}$ is stable, i.e., $\omega_{k}<0$ for all $k$ and hence $\omega^{*}=\max_{k}\{\omega_{k}\}<0$. Suppose then that $\phi$ increases or decreases monotonically. Suppose some $\omega_{k*}$ changes its sign from negative to positive at some $\phi^{*}$. Then, after that point, deviation towards the $\bm{z}_{k^{*}}$-direction improves the utility of the movers, and $\bar{\bm{x}}$ becomes unstable. In this sense, the corresponding eigenvector $\bm{z}_{k^{*}}$ represents the critical deviation pattern. ∎
@@ D $k=8$
These properties generalize to the $N$-region case. [^18] Akamatsu-Takayama-Ikeda-JEDC2012, Lemma 4.2, provides the analytical formulae for $\{\Theta_{k}\}$ and $\{\bm{z}_{k}\}$, while the interpretation of $\{\Theta_{k}\}$ as proximity gains is newly given in the present study. Lemma 5 in Appendix A reproduces the relevant part of the aforementioned lemma, adapted to our context. All $\{\Theta_{k}\}$ are positive and decrease in $\phi$. Each $\bm{z}_{k}$ corresponds to a deviation pattern with $k$ symmetric peaks as illustrated in Fig.˜7. Most importantly, the maximum and minimum proximity gains are unambiguously determined: for each $\phi$,
@@ A.1  Proof of Lemma˜1
First, we can represent $\Delta$ in terms of marginal migration from region $2$ to $1$: $\Delta(\epsilon)\coloneq\Delta(\bar{\bm{x}}+\frac{\epsilon}{2}(1,-1))=v_{1}(\bar{x}+\frac{\epsilon}{2},\bar{x}-\frac{\epsilon}{2})-v_{2}(\bar{x}+\frac{\epsilon}{2},\bar{x}-\frac{\epsilon}{2})$. Then, $\omega=\frac{\bar{x}}{\bar{v}}\Delta^{\prime}(\epsilon)|_{\epsilon=0}$, meaning that $\omega$ is proportional to the directional derivative of $\Delta$ with respect to the unit migration from region $2$ to $1$, as $\frac{1}{2}(1,-1)$ is a normalized vector. Here, we used the fact that $\frac{\partial{v_{1}}}{\partial{x_{1}}}(\bar{\bm{x}})=\frac{\partial{v_{2}}}{\partial{x_{2}}}(\bar{\bm{x}})$ and $\frac{\partial{v_{1}}}{\partial{x_{2}}}(\bar{\bm{x}})=\frac{\partial{v_{2}}}{\partial{x_{1}}}(\bar{\bm{x}})$ due to the symmetry of the regions at $\bar{\bm{x}}$. The stability condition based on the sign of $\omega$ is valid for general $\bm{v}$ provided that $\bar{v}>0$.
@@ A.3  Proof of Proposition˜1
Thus, we only need the eigenvalues of $\mathbf{V}$. A useful fact is that, if $\{(\Theta_{k},\bm{z}_{k})\}$ are the eigenpairs (eigenvalue–eigenvector pairs) of the normalized proximity matrix $\mathbf{D}$, then the eigenpairs of $\mathbf{V}=\Omega(\mathbf{D})$ are given by $\{(\Omega(\Theta_{k}),\bm{z}_{k})\}$ (e.g., Horn-Johnson-Book2012, Section 1.1). Thus, we need the eigenpairs of $\mathbf{D}$. For the sake of simplicity, we assume that $N$ is a multiple of four. Then, we have the following lemma.
@@
Assume Assumption˜C. The largest eigenvalue of $\mathbf{D}$ is $\Theta_{0}\equiv 1$, with $\bm{z}_{0}=(1,1,\ldots,1)$ being the associated eigenvector. Including $\Theta_{0}$, there are $\frac{N}{2}+1$ distinct eigenvalues. Every eigenvalue $\Theta_{k}$ ($k\neq 0$) is a strictly decreasing function of $\phi$, with $\lim_{\phi\downarrow 0}\Theta_{k}=1$ and $\lim_{\phi\uparrow 1}\Theta_{k}=0$. Let $\Theta_{\text{max}}$ denote the largest eigenvalue and $\Theta_{\text{min}}$ denote the smallest eigenvalue of $\mathbf{D}$ excluding $\Theta_{0}$, respectively. Further, assume that $N$ is a multiple of four. Then,
@@
By definition, there is $\Theta^{*}$ such that $\Omega^{\sharp}(\Theta)<0$ for all $\Theta\in(\Theta^{*},1)$, that $\Omega^{\sharp}(\Theta^{*})=0$, and that $\Omega^{\sharp}(\Theta)>0$ for all $\Theta\in(0,\Theta^{*})$. By Lemma˜5, $\{\Theta_{k}(\phi)\}$ are strictly decreasing from $1$. Thus, $\bar{\bm{x}}$ is stable if and only if $\Theta_{k}\in(\Theta^{*},1)$, so that $\omega^{\sharp}_{k}=\Omega^{\sharp}(\Theta_{k})<0$, for all $k$, i.e., if $\Theta^{*}<\min_{k}\Theta_{k}=\Theta_{N/2}$. Thus, $\bar{\bm{x}}$ is stable for all $(0,\phi^{*})$ where $\phi^{*}=(1-\sqrt{\Theta^{*}})/(1+\sqrt{\Theta^{*}})$ is the unique solution for $\Theta_{N/2}=\Theta^{*}$. Because $\Omega^{\sharp}(\Theta)>0$ for all $\Theta\in(0,\Theta^{*})$ and $\Theta_{N/2}$ is strictly decreasing, $\bar{\bm{x}}$ is unstable for all $(\phi^{*},1)$ because $\omega_{N/2}^{\sharp}>0$ for the range.
@@
In Lemma˜5, we assume that $N$ is a multiple of four to ensure $\min_{k}\{\Theta_{k}\}=\Theta_{\frac{N}{2}}$. This is inconsequential for the broad implication of Proposition˜1 on spatial patterns. If $N$ is an even, $\min_{k}\{\Theta_{k}\}=\min\{\Theta_{{\frac{N}{2}}-1},\Theta_{\frac{N}{2}}\}$. If $N$ is an odd, $\min_{k}\{\Theta_{k}\}=\min\{\Theta_{\lfloor\frac{N}{2}\rfloor},\Theta_{\lfloor\frac{N}{2}\rfloor-1}\}$. Thus, $\min_{k}\{\Theta_{k}\}$ corresponds to a polycentric direction, except for the case $N=2$ or $3$ in which polycentric patterns cannot occur. ∎
@@
The other prediction is that, if $M$ same-sized agglomerations are equidistantly placed on a circle, a symmetry-breaking bifurcation may reduce their number to $K<M$, with $K$ again dividing $N$ and the agglomerations remaining equidistant.
Proposition˜1 (a) implies that Type G models yield $\frac{N}{2}$ agglomerations, with further bifurcations of the form $\frac{N}{2}\to\frac{N}{4}\to\frac{N}{8}\to\cdots\to 2\to 1$ expected if $N$ is a power of two (Akamatsu-Takayama-Ikeda-JEDC2012; ikeda2012spatial; Osawa-et-al-JRS2017). For Type L, Takayama-Ikeda-Thisse-RSUE2020 confirmed the emergence of single-peaked or monocentric patterns in the (Murata-Thisse-JUE2005) model. AMT-2016 formally compares Forslid-Ottaviano-JoEG2003 (Type G) and Helpman-Book1998 (Type L). All available formal results in the literature corroborates with Proposition˜1 and the numerical examples in this study. ∎
@@ Appendix C Eight regions
Type L model. Figure˜C.3 considers a Type L model by Allen-Arkolakis-QJE2014 (Section˜F.2.4; we set $\alpha=0.5$, $\beta=-0.3$, and $\sigma=6.0$). The model incorporates a local dispersion force but no global dispersion force. The uniform equilibrium $\bar{\bm{x}}$ is stable when transport costs are low (when $\phi$ is close to $1$). If we start from $\bar{\bm{x}}$ and consider the process of a monotonic decrease in $\phi$ from $\phi\approx 1$, then a unimodal pattern emerges due to the bifurcation at $\phi^{**}$ [Proposition˜1 (a)]. This is the bifurcation in the model. When $\phi$ decreases further, the spatial pattern smoothly converges to a full concentration in a single region in the lower extreme ($\phi\approx 0$). The local dispersion force is less important than the benefits of agglomeration when interregional transportation is prohibitively costly. Mobile agents prefer concentrating on a smaller number of regions because of the agglomeration forces. As $\phi$ increases, agglomeration force due to costly transportation diminishes, and the relative rise in the local dispersion force induces a crowding-out from the populated region to the adjacent regions. As a result, the spatial pattern gradually flattens and connects to $\bar{\bm{x}}$ at $\phi^{**}$. We can interpret the region at the mode of population distribution (region $i$ such that $x_{i}>x_{i-1}$ and $x_{i}>x_{i+1}$ where mod $N$ for indices) as the location of an agglomeration. Then, this model endogenously produces at most one agglomeration.
@@
Type LG model. With both local and global dispersion forces, Type LG models exhibit an interplay between the number of agglomerations, spacing between them (as in Type G models), and the spatial extent of each agglomeration (as in Type L models). Figure˜2A shows the evolution of the number of agglomerations in the course of increasing $\phi$ under the Pfluger-Suedekum-JUE2008’s model (Section˜F.2.3; we set $\mu=0.4$, $\sigma=2.5$, $L=4$, $\gamma=0.5$, and $a_{i}=1$). The number of agglomerations in a spatial distribution is defined by that of the local maxima therein. Figure˜2A exhibits the mixed characteristics of Figs.˜C.3 and C.3, as expected. When $\phi<\phi^{*}$ or $\phi>\phi^{**}$, $\bar{\bm{x}}$ is stable. We interpret the number of agglomerations in $\bar{\bm{x}}$ as either $8$ (for a low $\phi$) or $1$ (for a high $\phi$) to acknowledge that $\bar{\bm{x}}$ at the low and high levels of $\phi$ are distinct. When $\phi$ gradually increases from $\phi\approx 0$, the number of agglomerations reduces from $8\to 4\to 2\to 1$ as in the Type G models (Fig.˜C.3), whereas it is always $1$ in the latter stage as per the Type L models (Fig.˜C.3). The initial stage is governed by a decline in the global dispersion force, while the later stage is marked by a relative rise of the local dispersion force.

### v9

@@ 2.1  The Beckmann model: A “local” dispersion force
While $\bar{\bm{x}}$ is an equilibrium, with both agglomeration and dispersion forces, $\bar{\bm{x}}$ is not stable if the former dominates the latter. To assess the stability of $\bar{\bm{x}}$, we examine the incentive of movers. Consider a small mass of agents moving from region 2 to 1, so that $x_{1}$ rises and $x_{2}$ falls symmetrically. If $\Delta$ decreases under this perturbation, then $\Delta<0$ after the move: utility in region 2 exceeds that in region 1, and the movers will prefer to return. This induces a restoring force and hence $\bar{\bm{x}}$ is locally stable. If $\Delta$ increases, then $\Delta>0$ after the move: the movers prefer to stay in region 1, and the deviation attracts additional movers now that the utility is higher in region 11. Hence, $\bar{\bm{x}}$ is unstable.
@@ 2.2  The Braid model: A “global” dispersion force
Here, $y_{i}(\bm{x})$ is the income of an agent in the region ii. Each region is endowed with one unit of fixed expenditure, which is allocated to all agents according to accessibility. Suppose that the share received by an agent in region $i$ from region $m$ is given by [^7] Note that the total expenditure of the region $m$ is equal to one: $\sum_{i\in\mathcal{I}}x_{i}s_{i\mid m}=1$. Possible microfoundations for this toy model are abstracted away for brevity.
@@
demonstrating that a marginal increase in agents in any region $j\in\{1,2\}$ has negative impacts on $s_{i\mid m}$. This is simply because it increases the denominator of $s_{i\mid m}$. Whether income $y_{i}=\sum_{j}s_{i\mid j}$ as a whole increases or not after a migration shock depends on the relative magnitudes of these impacts. Specifically, analogous to the utility gain $\omega$, we can compute the income gain of a marginal mover from region 22 to 11: [^8] From Eq. 9, $\frac{\partial{y_{1}}}{\partial{x_{1}}}=-(s_{1\mid 1})^{2}-(s_{1\mid 2})^{2}=-\frac{1+\phi^{2}}{\bar{x}^{2}(1+\phi)^{2}}$ and $\frac{\partial{y_{2}}}{\partial{x_{1}}}=-s_{2\mid 1}\cdot s_{1\mid 1}-s_{2\mid 2}\cdot s_{1\mid 2}=-\frac{2\phi}{\bar{x}^{2}(1+\phi)^{2}}$.
@@ 3  Many regions
This section examines how the proposed taxonomy of spatial models maps to the endogenous spatial patterns and their comparative statics in an NN-region economy. All variables and functions (e.g., $\bm{x}$, $\bm{v}$, $\phi_{ij}]$, $\mathbf{D}$) are straightforwardly extended. The set of regions is now denoted by $\mathcal{I}\equiv\{1,2,\ldots,N\}$.
@@
The proximity matrix is given by $\phi_{ij}=\phi^{\ell_{ij}}$, where $\phi\in(0,1)$ is the ease of transport between two consecutive regions, and $\ell_{ij}\equiv\min\{|i-j|,N-|i-j|\}$ is the distance between regions $i$ and $j$ over the circumference. All regions are symmetric regarding their local fundamentals (e.g., innate amenity or productivity). In addition, NN is a multiple of four. [^14] This restriction on NN is only for expositional simplicity. See Remark 3 in Appendix A.
@@ 3.1  The stability of the symmetric equilibrium
Endogenous agglomeration occurs when the symmetric equilibrium $\bar{\bm{x}}$ is unstable, that is, when it does not withstand small migration shocks. In an economy with NN regions, such shocks can be represented as
@@
where $\epsilon$ is a sufficiently small scalar. The vector $\bm{z}=(z_{i})_{i\in\mathcal{I}}$ is a deviation pattern: $z_{i}>0$ indicates an inflow into region $i$ and $z_{i}<0$ an outflow. Since the total population is fixed, we require $\sum_{i\in\mathcal{I}}z_{i}=0$. The NN-region setting admits a substantially richer set of deviation patterns than the two-region case, because $(\bar{x}+\epsilon,\bar{x}-\epsilon)$ is the only possible perturbation in the two-region setting. Below, we normalize $\|\bm{z}\|=1$.
@@
In the Beckmann model, we again have $\mathbf{V}=-\beta\mathbf{I}+\alpha\mathbf{D}$, where $\mathbf{I}$ and $\mathbf{D}$ are replaced by the NN-region identity and row-normalized proximity matrices. [^17] Observe that Eq. 2 is defined for general NN. The $k$th eigenvalue of $\mathbf{V}$ is then $\omega_{k}=-\beta+\alpha\Theta_{k}$ because $\mathbf{D}\bm{z}_{k}=\Theta_{k}\bm{z}_{k}$ implies $\mathbf{V}\bm{z}_{k}=(-\beta+\alpha\Theta_{k})\bm{z}_{k}$. The stability of $\bar{\bm{x}}$ is determined by the sign of $\omega^{*}=\max_{k}\{\omega_{k}\}$. Consider a value of $\phi$ such that $\bar{\bm{x}}$ is stable, i.e., $\omega_{k}<0$ for all $k$ and hence $\omega^{*}=\max_{k}\{\omega_{k}\}<0$. Suppose then that $\phi$ increases or decreases monotonically. Suppose some $\omega_{k*}$ changes its sign from negative to positive at some $\phi^{*}$. Then, after that point, deviation towards the $\bm{z}_{k^{*}}$-direction improves the utility of the movers, and $\bar{\bm{x}}$ becomes unstable. In this sense, the corresponding eigenvector $\bm{z}_{k^{*}}$ represents the critical deviation pattern. ∎
@@ D $k=8$
These properties generalize to the NN-region case. [^18] Akamatsu-Takayama-Ikeda-JEDC2012, Lemma 4.2, provides the analytical formulae for $\{\Theta_{k}\}$ and $\{\bm{z}_{k}\}$, while the interpretation of $\{\Theta_{k}\}$ as proximity gains is newly given in the present study. Lemma 5 in Appendix A reproduces the relevant part of the aforementioned lemma, adapted to our context. All $\{\Theta_{k}\}$ are positive and decrease in $\phi$. Each $\bm{z}_{k}$ corresponds to a deviation pattern with $k$ symmetric peaks as illustrated in Fig.˜7. Most importantly, the maximum and minimum proximity gains are unambiguously determined: for each $\phi$,
@@ A.1  Proof of Lemma˜1
First, we can represent $\Delta$ in terms of marginal migration from region 22 to 11: $\Delta(\epsilon)\coloneq\Delta(\bar{\bm{x}}+\frac{\epsilon}{2}(1,-1))=v_{1}(\bar{x}+\frac{\epsilon}{2},\bar{x}-\frac{\epsilon}{2})-v_{2}(\bar{x}+\frac{\epsilon}{2},\bar{x}-\frac{\epsilon}{2})$. Then, $\omega=\frac{\bar{x}}{\bar{v}}\Delta^{\prime}(\epsilon)|_{\epsilon=0}$, meaning that $\omega$ is proportional to the directional derivative of $\Delta$ with respect to the unit migration from region 22 to 11, as $\frac{1}{2}(1,-1)$ is a normalized vector. Here, we used the fact that $\frac{\partial{v_{1}}}{\partial{x_{1}}}(\bar{\bm{x}})=\frac{\partial{v_{2}}}{\partial{x_{2}}}(\bar{\bm{x}})$ and $\frac{\partial{v_{1}}}{\partial{x_{2}}}(\bar{\bm{x}})=\frac{\partial{v_{2}}}{\partial{x_{1}}}(\bar{\bm{x}})$ due to the symmetry of the regions at $\bar{\bm{x}}$. The stability condition based on the sign of $\omega$ is valid for general $\bm{v}$ provided that $\bar{v}>0$.
@@ A.3  Proof of Proposition˜1
Thus, we only need the eigenvalues of $\mathbf{V}$. A useful fact is that, if $\{(\Theta_{k},\bm{z}_{k})\}$ are the eigenpairs (eigenvalue–eigenvector pairs) of the normalized proximity matrix $\mathbf{D}$, then the eigenpairs of $\mathbf{V}=\Omega(\mathbf{D})$ are given by $\{(\Omega(\Theta_{k}),\bm{z}_{k})\}$ (e.g., Horn-Johnson-Book2012, Section 1.1). Thus, we need the eigenpairs of $\mathbf{D}$. For the sake of simplicity, we assume that NN is a multiple of four. Then, we have the following lemma.
@@
Assume Assumption˜C. The largest eigenvalue of $\mathbf{D}$ is $\Theta_{0}\equiv 1$, with $\bm{z}_{0}=(1,1,\ldots,1)$ being the associated eigenvector. Including $\Theta_{0}$, there are $\frac{N}{2}+1$ distinct eigenvalues. Every eigenvalue $\Theta_{k}$ ($k\neq 0$) is a strictly decreasing function of $\phi$, with $\lim_{\phi\downarrow 0}\Theta_{k}=1$ and $\lim_{\phi\uparrow 1}\Theta_{k}=0$. Let $\Theta_{\text{max}}$ denote the largest eigenvalue and $\Theta_{\text{min}}$ denote the smallest eigenvalue of $\mathbf{D}$ excluding $\Theta_{0}$, respectively. Further, assume that NN is a multiple of four. Then,
@@
By definition, there is $\Theta^{*}$ such that $\Omega^{\sharp}(\Theta)<0$ for all $\Theta\in(\Theta^{*},1)$, that $\Omega^{\sharp}(\Theta^{*})=0$, and that $\Omega^{\sharp}(\Theta)>0$ for all $\Theta\in(0,\Theta^{*})$. By Lemma˜5, $\{\Theta_{k}(\phi)\}$ are strictly decreasing from 11. Thus, $\bar{\bm{x}}$ is stable if and only if $\Theta_{k}\in(\Theta^{*},1)$, so that $\omega^{\sharp}_{k}=\Omega^{\sharp}(\Theta_{k})<0$, for all $k$, i.e., if $\Theta^{*}<\min_{k}\Theta_{k}=\Theta_{N/2}$. Thus, $\bar{\bm{x}}$ is stable for all $(0,\phi^{*})$ where $\phi^{*}=(1-\sqrt{\Theta^{*}})/(1+\sqrt{\Theta^{*}})$ is the unique solution for $\Theta_{N/2}=\Theta^{*}$. Because $\Omega^{\sharp}(\Theta)>0$ for all $\Theta\in(0,\Theta^{*})$ and $\Theta_{N/2}$ is strictly decreasing, $\bar{\bm{x}}$ is unstable for all $(\phi^{*},1)$ because $\omega_{N/2}^{\sharp}>0$ for the range.
@@
In Lemma˜5, we assume that NN is a multiple of four to ensure $\min_{k}\{\Theta_{k}\}=\Theta_{\frac{N}{2}}$. This is inconsequential for the broad implication of Proposition˜1 on spatial patterns. If NN is an even, $\min_{k}\{\Theta_{k}\}=\min\{\Theta_{{\frac{N}{2}}-1},\Theta_{\frac{N}{2}}\}$. If NN is an odd, $\min_{k}\{\Theta_{k}\}=\min\{\Theta_{\lfloor\frac{N}{2}\rfloor},\Theta_{\lfloor\frac{N}{2}\rfloor-1}\}$. Thus, $\min_{k}\{\Theta_{k}\}$ corresponds to a polycentric direction, except for the case $N=2$ or 33 in which polycentric patterns cannot occur. ∎
@@
The other prediction is that, if $M$ same-sized agglomerations are equidistantly placed on a circle, a symmetry-breaking bifurcation may reduce their number to $K<M$, with $K$ again dividing NN and the agglomerations remaining equidistant.
Proposition˜1 (a) implies that Type G models yield $\frac{N}{2}$ agglomerations, with further bifurcations of the form $\frac{N}{2}\to\frac{N}{4}\to\frac{N}{8}\to\cdots\to 2\to 1$ expected if NN is a power of two (Akamatsu-Takayama-Ikeda-JEDC2012; ikeda2012spatial; Osawa-et-al-JRS2017). For Type L, Takayama-Ikeda-Thisse-RSUE2020 confirmed the emergence of single-peaked or monocentric patterns in the (Murata-Thisse-JUE2005) model. AMT-2016 formally compares Forslid-Ottaviano-JoEG2003 (Type G) and Helpman-Book1998 (Type L). All available formal results in the literature corroborates with Proposition˜1 and the numerical examples in this study. ∎
@@ Appendix C Eight regions
Type L model. Figure˜C.3 considers a Type L model by Allen-Arkolakis-QJE2014 (Section˜F.2.4; we set $\alpha=0.5$, $\beta=-0.3$, and $\sigma=6.0$). The model incorporates a local dispersion force but no global dispersion force. The uniform equilibrium $\bar{\bm{x}}$ is stable when transport costs are low (when $\phi$ is close to 11). If we start from $\bar{\bm{x}}$ and consider the process of a monotonic decrease in $\phi$ from $\phi\approx 1$, then a unimodal pattern emerges due to the bifurcation at $\phi^{**}$ [Proposition˜1 (a)]. This is the bifurcation in the model. When $\phi$ decreases further, the spatial pattern smoothly converges to a full concentration in a single region in the lower extreme ($\phi\approx 0$). The local dispersion force is less important than the benefits of agglomeration when interregional transportation is prohibitively costly. Mobile agents prefer concentrating on a smaller number of regions because of the agglomeration forces. As $\phi$ increases, agglomeration force due to costly transportation diminishes, and the relative rise in the local dispersion force induces a crowding-out from the populated region to the adjacent regions. As a result, the spatial pattern gradually flattens and connects to $\bar{\bm{x}}$ at $\phi^{**}$. We can interpret the region at the mode of population distribution (region $i$ such that $x_{i}>x_{i-1}$ and $x_{i}>x_{i+1}$ where mod NN for indices) as the location of an agglomeration. Then, this model endogenously produces at most one agglomeration.
@@
Type LG model. With both local and global dispersion forces, Type LG models exhibit an interplay between the number of agglomerations, spacing between them (as in Type G models), and the spatial extent of each agglomeration (as in Type L models). Figure˜2A shows the evolution of the number of agglomerations in the course of increasing $\phi$ under the Pfluger-Suedekum-JUE2008’s model (Section˜F.2.3; we set $\mu=0.4$, $\sigma=2.5$, $L=4$, $\gamma=0.5$, and $a_{i}=1$). The number of agglomerations in a spatial distribution is defined by that of the local maxima therein. Figure˜2A exhibits the mixed characteristics of Figs.˜C.3 and C.3, as expected. When $\phi<\phi^{*}$ or $\phi>\phi^{**}$, $\bar{\bm{x}}$ is stable. We interpret the number of agglomerations in $\bar{\bm{x}}$ as either 88 (for a low $\phi$) or 11 (for a high $\phi$) to acknowledge that $\bar{\bm{x}}$ at the low and high levels of $\phi$ are distinct. When $\phi$ gradually increases from $\phi\approx 0$, the number of agglomerations reduces from $8\to 4\to 2\to 1$ as in the Type G models (Fig.˜C.3), whereas it is always 11 in the latter stage as per the Type L models (Fig.˜C.3). The initial stage is governed by a decline in the global dispersion force, while the later stage is marked by a relative rise of the local dispersion force.